Part B-Practice the steps for a one-step problem Once youve identfed the initial and Sinal condtions, youre ready to solve
Answer #1

according to Boyle's law P1V1 = P2V2

initial volume V1 = 300 L

final volume V2 = 2.00 L

initial pressure P1 = 1.00 atm

so final pressure P2 = P1V1/V2= 1.00*300/2.00 = 150 atm

so the pressure is 150.00 atm